OpenStudy (mrnebula):

Can someone show me how to justify the steps in this? 11x-12+8x+2=180 19x-10+180 19x+190 x+10 11*10-12= 98 4*10+1=41 41*2=82 98+82=180 98+41=139 AEC=139

OpenStudy (sunshinehaggard):

i can

OpenStudy (sunshinehaggard):

you add the 11x and 8x together

OpenStudy (sunshinehaggard):

then the -12 +2 which gives you -10

OpenStudy (sunshinehaggard):

you add the 10 to the other side with the 180 and you get 190

OpenStudy (sunshinehaggard):

then you divide by 19x to the 180

OpenStudy (sunshinehaggard):

thats the first part i dont think i understand the others let me look

OpenStudy (sunshinehaggard):

11*10=110 then you subtract the 12 and you get 98

OpenStudy (sunshinehaggard):

4*10=40+1=41

OpenStudy (sunshinehaggard):

41*2=82

OpenStudy (sunshinehaggard):

you add the 98 and the 82 together to get 180

OpenStudy (sunshinehaggard):

then the 98 and 41 to receive the 139